OpenStudy (anonymous):

Precalc, please help? Find the parametric equations for the line passing through points at (-1/3, 1, 1) and (0, 5, -8). (Hint: equations are needed for x, y, and z) The answer is: x=-1/3+1/3t, y=1+4t, z=1-9t How do I find this answer?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I always think of these as P = P0 + t*V where P0 is the starting point, and V is the direction to find V, subtract the two points they give you. we have choices here. They used (-1/3, 1, 1) as the starting point

OpenStudy (anonymous):

(0, 5, 8)-(-1/3, 1, 1)=(1/3, 4, 7) So I can see how to get the equations for x and y, but I don't understand how to find the equation for z.

OpenStudy (phi):

if you use the numbers they gave you, you would have better luck (0, 5, -8) <-- notice -8

OpenStudy (anonymous):

Ahh I see I missed the negative. Haha thank you!